Suppose that the following five portfolios are lying on the efficient frontier of an opportunity set,determine the best (optimal)portfolio,if the investor's risk tolerance is 40%

Portfolio A B C D E
expected return 7.60 9.13 9.43 9.73 9.85
expected variance 0.19 0.52 0.61 0.74 0.75
standard deviation 4.32 7.24 7.79 8.59 8.67
Answer #1

Stocks should be compared on the basis of Co-efficient of variation. Co -efficient of variation calculates the risk per unit of return. So, the stock having least co-efficient of variation will be preferred.

Co-efficient of Variation = Standard deviation/ expected return

A B C D E
4.32/7.6 =0.568421 7.24/9.13= 0.79299 7.79/9.43= 0.826087 8.59/9.73=0.882837 8.67/9.85=0.880203

So, on the basis of co-efficient of variation, Stock A is the best stock